OpenStudy (anonymous):

A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Explain

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

that is what I was thinking

OpenStudy (anonymous):

7^2+10^2=15^2 When you use the Pythagoras they don't relate so it wouldn't be a right triangle

OpenStudy (anonymous):

Now it's either going to be acute or obtuse and for it to be acute the hypotenuse of a right triangle would be longer than the hypotenuse which is 15 and since it's not then it's an obtuse triangle.

OpenStudy (anonymous):

Yea that one was pretty easy I remembered from a couple of questions we did for the exams

OpenStudy (anonymous):

cool